# On solving basic equations over the semiring of functional digraphs

### v3

@@ 2      $g\leftarrow\gcd(p,q)$;
Let $p$ and $q$ be two positive integers such that $p|q$. Let $p_{1}^{h_{1}}\cdot\ldots\cdot p_{r}^{h_{r}}$ and $q_{1}^{k_{1}}\cdot\ldots\cdot q_{s}^{k_{s}}$ be the prime number factorizations of $p$ and $q$, respectively. Let us also define $\operatorname{factors}\left(p\right)=\left\{p_{1},\ldots,p_{r}\right\}$ and $\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$. Finally, let $\mathbf{E}$ be the set of all $p_{i}=q_{j}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$ such that $h_{i}<k_{j}$. Then, Algorithm 2 computes ${\Pi}_{\mathbf{E}}=\prod_{\mathbf{e}\in\mathbf{E}}\mathbf{e}$ without using the factorizations of $p$ and $q$. The time complexity of Algorithm 2 is $O(\log p\cdot\log\log p)$.
@@
By executing Lines $2$ and $3$ of Algorithm 2, we save in $g$ the product of the factors $q^{k_{j}}_{j}$ such that $q_{j}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$. However, the exponents of the factors $q_{j}$ in $g$ are those of $q$ while to compute $\mathbf{\Pi_{E}}$ we need those of $p$. This is the purpose of the while loop (Lines $5$-$9$). Indeed, by recursively squaring $d$, it will make the exponents of the factors of $q$ selected in $g$ grow bigger than those in $p$ and hence the $\gcd$ of line $8$ will select the factors and the corresponding exponents that we were looking for. Concerning the complexity, it is enough to remark that the number of iterations of the while loop is $O(\log\log p)$ in the worst case and that each iteration costs $O(\log(\min(q/p,p)))$. ∎
@@ 8           $g^{\prime}\leftarrow\gcd(d,p)$;
Concerning the complexity, we have that the $\gcd$ at Line $2$ can be computed in $O(\log(\min(p,q)))$. By Lemmas 1 and 2, we know that computing $e$ takes $O(\max(\log p\cdot\log\log p,\log p\cdot\log q))$ that is $O(\log p\cdot\log q)$ since $p$ divides $q$. Finally, the check at Line $6$ has time complexity $O(\log(\min(e,n)))\subseteq O(\log p)$ since $e$ is a divisor of $p$. We conclude that the time complexity is $O(\log p\cdot\log q)$. ∎
@@ 1 Function DEEP(p, q, n)
      Output : $true$ if $C^{1}_{p}\odot X=C^{n}_{q}$ has solutions, $false$ otherwise
2      if $\gcd(p,q)\neq p$ then   // $p$ does not divide $q$
@@ 6                $e\leftarrow\Pi_{\mathbf{E}}(p,\frac{q}{{\Pi}_{\mathbf{F}}(p,q)})$;
$C^{1}_{8400}\odot X=C^{6000}_{8316000}$
where $p=8400$, $q=8316000$ and $n=6000$. Let us first see how by means of the presented algorithms that we can verify the conditions of Theorem 3 without knowing the factorisations of the numbers involved. Later, we will see how the method acts at the level of the factorisations.
@@
Let us therefore study ${\Pi}_{\mathbf{E}}(\frac{8316000}{11},8400)={\Pi}_{\mathbf{E}}(756000,8400)$. With $i=1$, $div=\frac{756000}{8400}=90$, $g=\gcd(90,8400)$ $=30$ and $g^{\prime}=\gcd(90^{2},8400)=\gcd(8100,8400)=300$. Since $g\neq g^{\prime}$, $i$ becomes $2$, $g$ takes value $300$ and $g^{\prime}$ becomes $\gcd(90^{3},8400)=\gcd(729000,8400)=600$. Hence, since $g$ and $g^{\prime}$ are still not equal, the method continues as follows.
@@
At this point, since $g=g^{\prime}$, the method returns $1200$. Finally, since $1200$ divides $n=6000$ we know that the equation admits of solution.
Let us now see what happens, from the point of view of the factorisations, by applying the previous method. Considering the values in input of this example, the factorisations are:
@@
The goal of ${\Pi}_{\mathbf{F}}$ is to compute the product of all $q_{j}^{k_{j}}$ such that $q_{j}\in\operatorname{factors}\left(q\right)\setminus\operatorname{factors}\left(p\right)$, and, in fact, the method calculates the following.
@@
Corollary 1.
Given four positive integers $p,q,n$ and $m$, the problem of deciding if the following equation
$C_{p}^{m}\cdot X=C_{q}^{n}$ (5)
has solutions has time complexity $O(\max(\log\min(m,n),\log p\cdot\log q))$.
Proof.
Assume that Equation (5) has a solution $\sum_{i=1}^{l}C_{t_{i}}^{s_{i}}$ then we have
$C^{n}_{q}=C_{p}^{m}\cdot\sum_{i=1}^{l}C_{t_{i}}^{s_{i}}=\sum_{i=1}^{l}C_{p}^{m}\cdot C_{t_{i}}^{s_{i}}=\sum_{i=1}^{l}C_{\operatorname{lcm}(p,t_{i})}^{ms_{i}\gcd(p,t_{i})}=C_{q}^{\sum_{i=1}^{l}\frac{mq}{p}\cdot\frac{s_{i}}{t_{i}}}=C_{q}^{\frac{mq}{p}\sum_{i=1}^{l}\frac{s_{i}}{t_{i}}}$
and hence Equation (5) has a solution iff $\frac{mq}{p}$ divides $n$ and $C_{p}^{1}\cdot X=C^{\frac{np}{mq}}_{q}$ has a solution. We conclude that the time complexity of deciding if Equation (5) has a solution is the max between $O(\log\min(\frac{mq}{p},n))$ (for checking if $\frac{mq}{p}$ divides $n$) and the time complexity of DEEP with parameters $p,q$ and $\frac{np}{mq}$ (see Theorem 4). ∎
Corollary 2.
Given $22s+$ 2 positive integers $p,m,q_{1},\ldots,q_{s},n_{1},\ldots,n_{s}$, the problem of deciding if the following equation
$C_{p}^{m}\cdot X=\sum_{i=1}^{s}C_{q_{i}}^{n_{i}}$ (6)
has solutions has time complexity $O(\max(\log\min(m,n),\log p\cdot\log q))$.
Proof.
It is clear that Equation (6) has solutions iff all the $s$ equations of type $C^{m}_{p}\cdot X=C^{n_{i}}_{q_{i}}$ have solutions. The last condition can be checked by applying a finite number of times Algorithm 3. Hence the time complexity is the same as the one of DEEP. ∎
@@
$\sum_{i=1}^{r}C_{p_{i}}^{m_{i}}\cdot X=C_{q}^{n}$ (7)
has solutions then $\sum_{k=1}^{r}m_{k}p_{k}$ must divide $nq$.
@@
Denote $d(q)$ the set of divisors of $q$. A generic solution to Equation (7) (if it exists) has the form $\hat{X}=\sum_{i=1}^{|d(q)|}C_{t_{i}}^{s_{i}}$, where $t_{i}$ is the $i$-th divisor of $q$ (divisors are considered ordered by the standard order on integers) and $s_{i}$ is an integer (possibly 00). We have
$C^{n}_{q}=\sum_{k=1}^{r}C_{p_{k}}^{m_{k}}\cdot\sum_{i=1}^{|d(q)|}C_{t_{i}}^{s_{i}}=\sum_{k=1}^{r}\sum_{i=1}^{|d(q)|}C_{p_{k}}^{m_{k}}\cdot C_{t_{i}}^{s_{i}}=\sum_{k=1}^{r}\sum_{i=1}^{|d(q)|}C_{\operatorname{lcm}(p_{k},t_{i})}^{m_{k}s_{i}\gcd(p_{k},t_{i})}$ (8)
@@
We deduce that if Equation (10) has solutions then $\sum_{k=1}^{r}m_{k}p_{k}$ divides $np$. ∎
Let us make some remarks concerning the proof of the previous results. Assume that $H=\frac{nq}{\sum_{k=1}^{r}m_{k}p_{k}}$ is an integer. Then, the Diophantine equation
@@
has solutions iff $\gcd(t_{1},\ldots,t_{|d(q)|})=q$ divides $H$, which is always true if $\sum_{k=1}^{r}m_{k}p_{k}$ divides $n$. With standard techniques (see for instance [7, Theorem 5.1, page 213]), one can recursively find the solutions in logarithmic time in $q$. We point out that it might happen that all the solutions are non admissible i.e., every solution contains at least one $s_{i}<0$ for $i\in\left\{1,\ldots,r\right\}$. This issue can also be settled in logarithmic time in $q$ (see the considerations following Theorem 5.1 in [7, pages 213-214]). However, the big problem that we have is that we need to compute the divisors of $q$ and, as far as we know, the best algorithms for doing that have time complexity which is a polynomial in $q$ (see [1] for details).
@@
Denote $d(q)$ the set of divisors of $q$. From Equation (8) in the proof of Proposition 2, it is easy to check that $\hat{X}=\sum_{i=1}^{|d(q)|}C_{t_{i}}^{s_{i}}$ is a solution if and only if
@@
Again, we would like to stress that solutions to Equation (13) can be found in logarithmic time in $\min(m_{1},\ldots,m_{n})$ giving raise to $r$ equations
@@
which can be also solved in logarithmic time provided that divisors $t_{i}$ of $q$ are known. Hence, once again the problem of finding solutions to Equation (7) can be solved in polynomial time in the value of $q$ (more precisely in $O(\max(q,\gcd(m_{1},\ldots,m_{k})))$). However, for some particular values of $q$, Equation (7) can be solved in an effective way. An example is shown by the following proposition.
@@
Given $r+1$ positive integers $p_{1},\ldots,p_{r},q$ and $r+2$ integers $m_{1},\ldots,m_{r},t$ and $n$. Let $R=\left\{1,\ldots,r\right\}$. Assume $q$ is prime and that for $i\in R$ we have $p_{i}\neq q^{t}$. Then, the following equation
$\sum_{i=1}^{r}C_{p_{i}}^{m_{i}}\cdot X=C^{n}_{q^{t}}$ (14)
has solutions iff $\sum_{i=1}^{r}m_{i}p_{i}$ divides $n$ and for $i\in R$, $p_{i}=q^{t_{i}}$ for some integer $t_{i}<t$. Moreover, if a solution exists, then it is unique.
@@
$\sum_{i=1}^{r}C_{p_{i}}^{m_{i}}\cdot\sum_{k=1}^{l}C_{v_{k}}^{u_{k}}=\sum_{i=1}^{r}\sum_{k=1}^{l}C_{p_{i}}^{m_{i}}\cdot C_{v_{k}}^{u_{k}}=\sum_{i=1}^{r}\sum_{k=1}^{l}C_{\operatorname{lcm}(p_{i},v_{k})}^{m_{i}u_{k}\gcd(p_{i},v_{k})}=C^{n}_{q^{t}}$
which implies $\operatorname{lcm}(p_{i},v_{k})=q^{t}$. Since $q$ is prime, we deduce that either $p_{i}=q^{t}$ or $v_{k}=q^{t}$. Using the hypothesis we have that $v_{k}=q^{t}$ for all $k\in\left\{1,\ldots,l\right\}$. Hence, the solution can be written as $C_{q^{t}}^{u}$. Now, for each $i\in R$ we have $\gcd(p_{i},q^{t})=p_{i}$. Hence, $C_{q^{t}}^{u}$ is a solution iff the following holds
$\sum_{i=1}^{r}C_{p_{i}}^{m_{i}}\cdot C_{q^{t}}^{u}=\sum_{i=1}^{r}C_{q^{t}}^{m_{i}up_{i}}=C^{\sum_{i=1}^{r}m_{i}up_{i}}_{q^{t}}=C^{n}_{q^{t}}$
@@
()($\Leftarrow$)( ⇐ ) It is not difficult to see that if $\sum_{i=1}^{r}m_{i}p_{i}$ divides $n$, then $C^{\frac{n}{\sum_{i=1}^{r}m_{i}p_{i}}}_{q^{t}}$ is a solution provided that $p_{i}$ divides $q^{t}$ which implies that $p_{i}=q^{t_{i}}$ for some $t_{i}\leq t$.
Concerning the uniqueness of the solution, it is enough to remark that the question is equivalent to ask for solutions of a Diophantine equation on a single variable. ∎
4 Conclusion
This paper proves that DEEP  and some of its variants are in ¶. For some other variants, we were able to prove only polynomially testable necessary conditions. These results will greatly improve the performances of the software pipeline for solving general polynomial equations over functional graphs proposed in [4] when one is interested in the existence of solutions and not to their enumeration.
Several improvements to our results are possible. The first one is, of course, to reproduce similar results for general functional graphs, taking into accounts new recent results from [5] and [8].

### v4

@@ 2      $g\leftarrow\gcd(p,q)$;
Let $p$ and $q$ be two positive integers such that $p|q$. Let $p_{1}^{h_{1}}\cdot\ldots\cdot p_{r}^{h_{r}}$ and $q_{1}^{k_{1}}\cdot\ldots\cdot q_{s}^{k_{s}}$ be the prime number factorizations of $p$ and $q$, respectively. Let us also define $\operatorname{factors}\left(p\right)=\left\{p_{1},\ldots,p_{r}\right\}$ and $\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$. Finally, let $\mathbf{E}$ be the set of all $p_{i}=q_{j}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$ such that $h_{i}<k_{j}$. Then, Algorithm 2 computes ${\Pi}_{\mathbf{E}}=\prod_{\mathbf{e}\in\mathbf{E}}\mathbf{e}$ without using the factorizations of $p$ and $q$. The worst-case time complexity of Algorithm 2 is $O(n^{2}\log n)$.
@@
By executing Lines $2$ and $3$ of Algorithm 2, we save in $g$ the product of the factors $q^{k_{j}}_{j}$ such that $q_{j}\in\operatorname{factors}\left(p\right)\cap\operatorname{factors}\left(q\right)$. However, the exponents of the factors $q_{j}$ in $g$ are those of $q$ while to compute $\mathbf{\Pi_{E}}$ we need those of $p$. This is the purpose of the while loop (Lines $5$-$9$). Indeed, by recursively squaring $d$, it will make the exponents of the factors of $q$ selected in $g$ grow bigger than those in $p$ and hence the $\gcd$ of line $8$ will select the factors and the corresponding exponents that we were looking for. Concerning the complexity, it is enough to remark that initialisation part of the algorithm (Lines $2$ to $4$) costs $O(n^{2})$ and that the number of iterations of the while loop is $O(\log n)$ in the worst-case with a cost of $O(n^{2})$ per each iteration. ∎
@@ 8           $g^{\prime}\leftarrow\gcd(d,p)$;
Concerning the complexity, we have that the $\gcd$ in Line $2$ can be computed in $O(s^{2})$. By Lemmata 1 and 2, we know that computing $e$ takes $O(\max(s^{3},s^{2}\log s,s^{2}))$ that is $O(s^{3})$. Remark that the size in bits of the quantities involved in the calculation of $e$ is bounded by $s$. Indeed, both $\Pi_{F}$ and $\Pi_{E}$ return a divisor of $q$.
Finally, the checks in Lines $2$ and $6$ have complexity $O(s^{2})$ since $e$ is a divisor of $p$ and, in its turn, $p$ is a divisor of $q$. We conclude that the time complexity is $O(s^{3})$. ∎
@@ 1 Function DEEP(p, q, n)
      Output : $true$ if $C_{p}\cdot X=nC_{q}$ has solutions, $false$ otherwise
2      if $p$ does not divide $q$ then
@@ 6                $e\leftarrow\Pi_{\mathbf{E}}(p,\frac{q}{{\Pi}_{\mathbf{F}}(p,q)})$;
$C_{8400}\cdot X=6000C_{8316000}$
where $p=8400$, $q=8316000$ and $n=6000$. Let us first see how by means of the presented algorithms that we can verify the conditions of Theorem 3 without knowing the factorizations of the numbers involved. Later, we will see how the method acts at the level of the factorizations.
@@
Let us therefore study ${\Pi}_{\mathbf{E}}(\frac{8316000}{11},8400)={\Pi}_{\mathbf{E}}(756000,8400)$. With $i=1$, $d=\frac{756000}{8400}=90$, $g=\gcd(90,8400)$ $=30$ and $g^{\prime}=\gcd(90^{2},8400)=\gcd(8100,8400)=300$. Since $g\neq g^{\prime}$, $i$ becomes $2$, $g$ takes value $300$ and $g^{\prime}$ becomes $\gcd(90^{3},8400)=\gcd(729000,8400)=600$. Hence, since $g$ and $g^{\prime}$ are still not equal, the method continues as follows.
@@
At this point, since $g=g^{\prime}$, the method returns $1200$. Finally, since $1200$ divides $n=6000$ we know that the equation admits a solution.
Let us now see what happens, from the point of view of the factorizations, by applying the previous method. Considering the values in input of this example, the factorizations are:
@@
so we have $\operatorname{factors}\left(p\right)=\left\{2,3,5,7\right\}$ and $\operatorname{factors}\left(q\right)=\left\{2,3,5,7,11\right\}$. The goal of ${\Pi}_{\mathbf{F}}$ is to compute the product of all $q_{j}^{k_{j}}$ such that $q_{j}\in\operatorname{factors}\left(q\right)\setminus\operatorname{factors}\left(p\right)$, and, in fact, the method calculates the following.
@@
The proof of the following proposition is trivial since it boils down to remark the the number of the vertices of the digraphs involved in the left part of the equation must equal the number of vertices of digraphs in the right part. However, we shall provide an alternative proof to explicit some relations that are discussed right after the end of the proof.
@@
$\sum_{i=1}^{r}m_{i}C_{p_{i}}\cdot X=nC_{q}$ (5)
has solutions then $\sum_{i=1}^{r}m_{i}p_{i}$ must divide $nq$.
@@
Denote $d(q)$ the set of divisors of $q$. A generic solution to Equation (5) (if it exists) has the form $\hat{X}=\sum_{i=1}^{|d(q)|}s_{i}C_{t_{i}}$, where $t_{i}$ is the $i$-th divisor of $q$ (divisors are considered ordered by the standard order on integers) and $s_{i}$ is an integer (possibly 00). We have
$nC_{q}=\sum_{k=1}^{r}m_{k}C_{p_{k}}\cdot\sum_{i=1}^{|d(q)|}s_{i}C_{t_{i}}=\sum_{k=1}^{r}\sum_{i=1}^{|d(q)|}m_{k}C_{p_{k}}\cdot s_{i}C_{t_{i}}=\sum_{k=1}^{r}\sum_{i=1}^{|d(q)|}m_{k}s_{i}\gcd(p_{k},t_{i})C_{\operatorname{lcm}(p_{k},t_{i})}$ (6)
@@
We deduce that if Equation (8) has solutions then $\sum_{k=1}^{r}m_{k}p_{k}$ divides $nq$. ∎
Let us make some remarks concerning the proof of the previous result. Assume that $H=\frac{nq}{\sum_{k=1}^{r}m_{k}p_{k}}$ is an integer. Then, the Diophantine equation
@@
has solutions iff $\gcd(t_{1},\ldots,t_{|d(q)|})$ divides $H$, which is always true if $\sum_{k=1}^{r}m_{k}p_{k}$ divides $n$. With standard techniques (see for instance [9, Theorem 5.1, page 213]), one can recursively find the solutions in linear time in the size of $q$ in bits. We point out that it might happen that all the solutions are non admissible i.e., every solution contains at least one $s_{i}<0$ for $i\in\left\{1,\ldots,r\right\}$. This issue can also be settled in linear time in the size of $q$ in bits (see the considerations following Theorem 5.1 in [9, pages 213-214]). However, the big problem that we have is that we need to compute the divisors of $q$ and, as far as we know, the best algorithms for computing them have worst-case time complexity which is sub-exponential in the size of $q$ (see [1] for details).
@@
Denote $d(q)$ the set of divisors of $q$. From Equation (6) in the proof of Proposition 2, it is easy to check that $\hat{X}=\sum_{i=1}^{|d(q)|}s_{i}C_{t_{i}}$ is a solution if and only if
@@
Again, we would like to stress that solutions to Equation (11) can be found in quadratic time in in the size of the input giving raise to $r$ equations
@@
which can be also solved in quadratic time provided that divisors $t_{i}$ of $q$ are known. Hence, once again the problem of finding solutions to Equation (5) can be solved in polynomial time in the size of $q$ in bits (of course, always knowning the divisors of $q$). However, for some particular values of $q$, Equation (5) can be solved in a more effective way. An example is shown by the following proposition.
@@
Given $r+1$ positive integers $p_{1},\ldots,p_{r},q$ and $r+2$ integers $m_{1},\ldots,m_{r},t$ and $n$. Let $R=\left\{1,\ldots,r\right\}$. Assume $q$ is prime and that for every $i\in R$ we have $p_{i}\neq q^{t}$. Then, the following equation
$\sum_{i=1}^{r}m_{i}C_{p_{i}}\cdot X=nC_{q^{t}}$ (12)
has solutions iff $\sum_{i=1}^{r}m_{i}p_{i}$ divides $n$ and for every $i\in R$, $p_{i}=q^{t_{i}}$ for some integer $t_{i}<t$. Moreover, if a solution exists, then it is unique.
@@
$\sum_{i=1}^{r}m_{i}C_{p_{i}}\cdot\sum_{k=1}^{l}u_{k}C_{v_{k}}=\sum_{i=1}^{r}\sum_{k=1}^{l}m_{i}C_{p_{i}}\cdot u_{k}C_{v_{k}}=\sum_{i=1}^{r}\sum_{k=1}^{l}m_{i}u_{k}\gcd(p_{i},v_{k})C_{\operatorname{lcm}(p_{i},v_{k})}=nC_{q^{t}}$
which implies $\operatorname{lcm}(p_{i},v_{k})=q^{t}$. Since $q$ is prime, we deduce that either $p_{i}=q^{t}$ or $v_{k}=q^{t}$. Using the hypothesis we have that $v_{k}=q^{t}$ for all $k\in\left\{1,\ldots,l\right\}$. Hence, the solution can be written as $uC_{q^{t}}$. Now, for every $i\in R$, we have $\gcd(p_{i},q^{t})=p_{i}$. Hence, $uC_{q^{t}}$ is a solution iff the following holds
$\sum_{i=1}^{r}m_{i}C_{p_{i}}\cdot uC_{q^{t}}=\sum_{i=1}^{r}m_{i}up_{i}C_{q^{t}}=\left(\sum_{i=1}^{r}m_{i}up_{i}\right)C_{q^{t}}=nC_{q^{t}}$
@@
()($\Leftarrow$)( ⇐ ) It is not difficult to see that if $\sum_{i=1}^{r}m_{i}p_{i}$ divides $n$, then $\frac{n}{\sum_{i=1}^{r}m_{i}p_{i}}C_{q^{t}}$ is a solution provided that $p_{i}$ divides $q^{t}$ which implies that $p_{i}=q^{t_{i}}$ for some $t_{i}\leq t$.
Concerning the uniqueness of the solution, it is enough to remark that the question is equivalent to asking for solutions of a Diophantine equation on a single variable. ∎
Clearly, the complexity of the algorithm verifying the conditions of Propositions 2, 3 and 4 is quadratic in the size of the input.
4 Conclusions
This paper is concerned with DEEP i.e., the problem of deciding if a basic equation on permutations has a solution or not. We show that DEEP has cubic complexity. However, it is clear that with more care in the choice of the implementation of some of the components of the decision algorithm, we could obtain a time complexity located between $O(n^{1+\epsilon}\log^{k}n)$ and $O(n^{3})$. Our purpose was essentially to prove that the problem could be solved in polynomial time with a polynomial of reasonable degree, but it is clear that when the instances have large sizes, like in some practical applications, then the shift towards more complex implementations of the components should be taken into account.
For some other variants of DEEP, we were able to prove only some necessary conditions which are testable in quadratic time. However, they have some interest of their own as they might also be of help in pruning the search space for software that aims to solve general polynomial equations on functional digraphs such as the software pipeline proposed in [4].
Several extensions to our results are possible. The most natural one consists of moving from functional digraphs of permutations to larger classes of functional digraphs taking into account the new recent results of [5, 6] and [8].